OpenStudy (anonymous):

the sum of two numbers is 25 and their difference is 19.find the numbers

OpenStudy (anonymous):

Their sum is 25 1)x + y = 25 Their difference is 19 2) x - y = 19 x = 19 + y plugging that in to the first equation: (19 + y )+ y = 25 19 + y + y = 25 19 + 2y = 25 2y = 6 y = 3 now just plug that into one of the equations: x + y = 25 x + 3 = 25 x = 22